A 2.6 kg object is tied to the end of a cord and whirled in a horizontal circle of radius 3.3 m completing one revolution in 7 seconds. Determine the velocity of the object to one decimal place.

1
Expert's answer
2022-01-19T06:00:07-0500
"v=\dfrac{2\pi R}{T}=\dfrac{2\pi·3.3}{7}=3.0\text{ m/s}."